OpenStudy (anonymous):

Please help it is so simple Find the 13th term of the arithmetic sequence $1200, $964.32, $728.64, …

OpenStudy (anonymous):

Why is people so scared of arithmetic? everytime I post this people leave?

OpenStudy (anonymous):

It goes down by 235.68 each time

OpenStudy (anonymous):

THat helps

OpenStudy (anonymous):

tn=t1+(n-1)d is the formula

jimthompson5910 (jim_thompson5910):

tn = t1 + (n-1)*d tn = 1200 + (n-1)*(-235.68) tn = 1200 - 235.68n + 235.68 tn = 1435.68 - 235.68n tn = -235.68n + 1435.68 So the nth term is tn = -235.68n + 1435.68 This is assuming n starts at n = 1

OpenStudy (anonymous):

Nice